OpenStudy (anonymous):

A firecracker shoots up from a hill 140 feet high with an initial speed of 100 feet per second. Using the formula H(t) = −16t2 + vt + s, approximately how long will it take the firecracker to hit the ground?

OpenStudy (anonymous):

ok i will use calculus based approach are you familiar with derivatives?

OpenStudy (anonymous):

not really

OpenStudy (anonymous):

i think it is from the calculus course isn't it?

OpenStudy (anonymous):

im only in algebra 1

OpenStudy (anonymous):

anyway taking derivative of H(t) \[H'(t)=-16*2(t)+v=-32t+v\]

OpenStudy (anonymous):

are u okay with this ?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

ok set it equal to zero \[-32t+v=0\] solve for t \[32t=v\] \[t=\frac{v}{32}\] but v=100 \[t=\frac{100}{32}=3.125 \]

OpenStudy (anonymous):

put t=3.125 in the original equation let e know what you got?

OpenStudy (anonymous):

okay hold on

OpenStudy (anonymous):

do not forget to use s=0 and v=100 along with t=3.125 in the equation !

OpenStudy (anonymous):

i dont think i did it right? but i got h(t)=156...

OpenStudy (anonymous):

you got it correct!